### v1

# Adversarial Alignment: Ensuring Value Consistency in Large Language Models for Sensitive Domains

Abstract
With the wide application of large language models (LLMs), the problems of bias and value inconsistency in sensitive domains have gradually emerged, especially in terms of race, society and politics. In this paper, we propose an adversarial alignment framework, which enhances the value consistency of the model in sensitive domains through continued pre-training, instruction fine-tuning and adversarial training. In adversarial training, we use the Attacker to generate controversial queries, the Actor to generate responses with value consistency, and the Critic to filter and ensure response quality. Furthermore, we train a Value-Consistent Large Language Model, VC-LLM, for sensitive domains, and construct a bilingual evaluation dataset in Chinese and English. The experimental results show that VC-LLM performs better than the existing mainstream models in both Chinese and English tests, verifying the effectiveness of the method.
Warning: This paper contains examples of LLMs that are offensive or harmful in nature.
keywords:
Large Language Models; Adversarial Alignment; Value Consistency
journal: Data Intelligence
1 Introduction
The rapid development of large language models (LLMs, such as GPT-4 [2], Llama-3 [14], Qwen2.5 [30], and GLM-4 [17]) has led to their widespread application across various downstream tasks [41, 39, 23, 11]. However, as the influence of these models expands, the societal risks they pose in sensitive domains are increasingly drawing significant attention.
Recent studies have shown that LLMs exhibit value biases related to race [1], occupation [26], society [15], and politics [5, 24, 32], which stem from inherent value biases in the training data [28]. Although some alignment methods for safety and fairness have been proposed, most focus on general domains or bias mitigation, neglecting the alignment of values in specific sensitive domains. Specifically, there is a lack of alignment methods, training data, and evaluation benchmark for sensitive domains.
To address these challenges, we propose an adversarial alignment framework, which incorporates continued pre-training, instruction fine-tuning, and adversarial training. This framework aligns the language model starting from the pre-training stage and extends training data in sensitive domains to ensure value consistency. The core of the framework lies in adversarial training, which consists of three components: the Attacker generates queries in sensitive domains, the Actor produces responses based on value alignment principles, and the Critic reviews the generated responses to ensure value alignment.
Our contributions can be summarized as follows:
  • 1.
    We propose an adversarial alignment framework to generate value alignment datasets in sensitive domains, which automatically produces high-quality and challenging alignment data.
  • 2.
    Based on this framework, we train VC-LLM, a large language model with value consistency in Chinese sensitive domains, and publicly release the constructed value alignment dataset  [^1] https://github.com/Ad-Align/VC-LLM.
  • 3.
    We construct a bilingual value alignment evaluation benchmark in Chinese and English, and conduct comparative experiments and analysis on mainstream LLMs and VC-LLM based on this benchmark.
2 Related Work
2.1 Value Bias in LLMs
Numerous studies have revealed that LLMs exhibit value-laden biases across domains such as religion, social expectations, and political ideology. For example, Abid et al. [1] find that GPT-3 demonstrates significant anti-Muslim bias, while Salecha et al. [33] show that LLMs tend to overrepresent overly positive personality traits when simulating human characteristics. Bernardelle et al. [7] report that models exhibit inherent leanings in political compass tests, which are difficult to mitigate through simple prompt engineering. Similarly, LLMs can display political favoritism and sycophantic behavior toward different political parties [27, 6]. In addition, LLMs often show reluctance to answer political questions [21], diverse mental values [25], and differences in moral and political stances [22, 37].
2.2 Value Alignment in LLMs
Recent studies have focused on the alignment of LLMs for safety [31, 42, 36, 34, 35]. For example, Bai et al. [4] combine preference modeling with human feedback-based reinforcement learning to achieve safe alignment. These methods primarily aim to enhance the model’s helpfulness and safety, or align it with globally recognized values. However, their application in sensitive domains remains limited. To address this challenge, some research has begun exploring alignment in specific domains. For instance, Liu et al. [24] use reinforcement learning to reduce political bias in LLM responses, Agiza et al. [3] optimize political bias through instruction fine-tuning and preference adjustment, while Borah et al. [8] propose self-reflection strategies to detect and mitigate hidden biases. Although progress has been made in sensitive domains alignment, these methods often focus on identifying and mitigating bias, without addressing the alignment of domain-specific values. We propose an adversarial alignment framework, aiming to ensure alignment of LLMs in sensitive domains.
Figure 1: Overview of the adversarial training framework. The framework begins with sensitive words (e.g., Taiwan sovereignty). The Attacker generates challenging queries (e.g., What is the truth behind the Taiwan independence movement?), the Actor produces value-aligned responses (e.g., emphasizing national unity), and the Critic filters for high-quality Q&R (Query & Response) pairs to construct training data.
3 Adversarial Value Alignment for Sensitive Domains
We propose an adversarial alignment framework to enhance the value alignment of LLMs in various sensitive domains. We firstly introduce the framework, followed by its core component: adversarial training. Finally, we apply this framework in the context of Chinese sensitive domains, resulting in the trained VC-LLM.
3.1 Training Framework
The adversarial alignment framework consists of three stages:
Continued Pre-training. To mitigate the bias caused by value misalignment in the original pre-training data, we perform continued pre-training using value-aligned data from sensitive domains.
Instruction Fine-Tuning. In this stage, we design two instruction formats to improve the model’s value alignment in sensitive domains: (1) Question-Answering Task: Given a question, the model is required to generate a value-aligned answer; (2) Statement-Rebuttal Task: Given a value-inconsistent statement, the model needs to identify value conflicts and produce a rebuttal that reflects aligned values.
Adversarial Training. We introduce adversarial training to augment the training data and enhance the model’s robustness.
3.2 Adversarial Training
We propose a collaborative adversarial training process consisting of the Attacker, Actor, and Critic, designed to expand the training data and enhance the model’s value alignment and safety, as shown in Figure 1. Using the Attacker and Actor to generate queries and responses, this approach mitigates the biases and consistency issues that may arise from using a single model. Additionally, the Critic ensures the quality of training data. The pseudocode is as follows:
Algorithm 1 Adversarial Alignment Dataset Generation
Input: Sensitive word list $S$ (e.g., [Taiwan, Zhenbao Island, ...]), Template library $T$ Output: Adversarial alignment dataset $D_{\text{adv}}$ (query-response pairs)
1: Initialize $D_{\text{adv}}\leftarrow\emptyset$
2:for each sensitive topic $s\in S$ do
3:  $prompt\leftarrow\text{fill\_template}(\text{sample}(T),s)$
4:  $raw\_query\leftarrow\text{{Attacker}.generate\_query}(prompt)$
5:  $(raw\_query,response)\leftarrow$ Actor.process_query(raw_query, prompt.type)
6:  if not Critic.check_failure(raw_query,responseraw\_query,response) then
7:   $D_{\text{adv}}\leftarrow D_{\text{adv}}\cup\{(raw\_query,response)\}$
8:  end if
9:end for
10:return $D_{\text{adv}}$
Attacker. The Attacker is used to generate questions and statements for adversarial training. Since fine-tuning can weaken a model’s safety alignment [29], we fine-tune LLM to enable it to generate queries for sensitive domains.
Actor. The Actor is used to generate value-aligned responses for adversarial training. It is a more capable language model with stronger alignment performance. For misleading or biased questions, it should produce truthful and values-consistent answers. For statements with value inconsistencies, it generates rebuttals that reflect consistent values.
Critic. The Critic is used to evaluate and filter the responses generated by the Actor during adversarial training to ensure response quality and value alignment. It is a large language model that serves as an automatic evaluator. Specifically, the Critic filters out responses that are evasive, value-misaligned, or incomplete.
The prompt templates for the Attacker, Actor, and Critic are shown in A.
3.3 VC-LLM
Figure 2: Prompts and Examples in Instruction Fine-Tuning.
We train VC-LLM in Chinese sensitive domains using this framework.
During the continued pre-training stage, we use Llama-3-Chinese-8B [12] as the base model, which has been continuously pre-trained on Chinese corpora to enhance its language understanding. The continued pre-training data is approximately 33.7GB, covering five sensitive domains: Sovereignty, Human Rights, Religion, Politics, and Ethnicity. The data sources include official news reports, information from provincial and municipal government websites, relevant national and local policies and regulations, meeting records, and published books.
During the instruction fine-tuning stage, we construct the instruction fine-tuning dataset using authoritative data from various sources. Some datasets, such as foreign ministry spokesperson remarks, are directly converted from their question-and-answer formats. For unannotated data, we use the GLM-4-9B-Chat [17] to generate instruction pairs based on it (prompts and examples shown in Figure 2). General data are also included to ensure broad capabilities.
During the adversarial value alignment stage, we use the Qwen2.5-7B-Instruct [30] as the Attacker, due to its small parameter size and fast inference speed, making it suitable for generating large volumes of data. We fine-tune the model using misaligned responses from the Safe RLHF dataset [13] to ensure the model can generate queries in sensitive domains. For the Actor, we use the Qwen2.5-72B-Instruct, which is selected for its strong instruction-following ability and good safety alignment, making it ideal for generating aligned responses in sensitive domains. The GLM-4-9B-Chat is chosen as the Critic, as it is from a different series than both the Attacker and Actor, ensuring objectivity in evaluation, while also benefiting from faster inference speed. In total, we generate 64,238 adversarial alignment samples.
Details of the training data and hyperparameters are provided in B.
Figure 3: Prompt used to evaluate the LLM’s responses, with a maximum score of 5.
Model Sov HR Rel Pol Eth Total Avg
Baichuan2-7B-Chat 244 (3.75) 94 (3.13) 69 (3.63) 140 (3.33) 56 (3.11) 603 3.47
Baichuan2-13B-Chat 262 (4.03) 106 (3.53) 74 (3.89) 156 (3.71) 66 (3.67) 664 3.82
GLM-4-9B-Chat 249 (3.83) 115 (3.83) 83 (4.37) 165 (3.93) 69 (3.83) 681 3.91
Yi-6B-Chat 242 (3.72) 98 (3.27) 70 (3.68) 132 (3.14) 57 (3.17) 599 3.44
internlm2.5-7b-chat 106 (1.63) 47 (1.57) 29 (1.53) 57 (1.36) 37 (2.06) 276 1.59
TigerBot-7B 119 (1.83) 39 (1.30) 34 (1.79) 81 (1.93) 37 (2.06) 310 1.78
Ziya2-13B-Chat 200 (3.08) 81 (2.70) 59 (3.11) 119 (2.83) 42 (2.33) 501 2.88
Llama-3-Ch-8B-v3 181 (2.78) 84 (2.80) 51 (2.68) 111 (2.64) 52 (2.89) 479 2.75
Qwen2.5-7B-Instruct 304 (4.68) 115 (3.83) 88 (4.63) 156 (3.71) 70 (3.89) 733 4.21
Qwen2.5-14B-Instruct 296 (4.55) 121 (4.03) 88 (4.63) 163 (3.88) 78 (4.33) 746 4.29
Qwen2.5-72B-Instruct 302 (4.65) 117 (3.90) 92 (4.84) 176 (4.19) 79 (4.39) 766 4.40
GPT-4o 202 (3.11) 86 (2.87) 48 (2.53) 123 (2.93) 39 (2.17) 498 2.86
VC-LLM 305 (4.69) 143 (4.77) 95 (5.00) 206 (4.90) 87 (4.83) 836 4.80
Table 1: Chinese evaluation results of VC-LLM and other baseline models. The total score is out of 870 points, with the breakdown as follows: Sovereignty (325 points), Human Rights (150 points), Religion (95 points), Politics (210 points), and Ethnicity (90 points). We use the form of total score (avg score) in the table.
4 Experiments and Analysis
4.1 Experimental Setups
We construct a bilingual value evaluation dataset (Chinese and English), covering five sensitive domains: Sovereignty (Sov), Human Rights (HR), Religion (Rel), Politics (Pol), and Ethnicity (Eth), further divided into 31 subcategories, with details in C.
We compared evaluations from human annotators with those from Qwen2.5-7B-Instruct and Qwen2.5-72B-Instruct, both using prompt-based assessment. The results demonstrate high consistency between the models and human judgments, with exact match rates of 87.36% (Zh) and 87.93% (En) (calculated as $\frac{1}{N}\sum_{i=1}^{N}1(s_{i}^{\text{pred}}=s_{i}^{\text{gold}})$) and tolerance-based match rates of 96.55% and 98.85% (calculated as $\frac{1}{N}\sum_{i=1}^{N}1(|s_{i}^{\text{pred}}-s_{i}^{\text{gold}}|\leq 1)$), respectively. Considering the faster inference speed of Qwen2.5-7B-Instruct, we selected it for evaluation (prompt in Figure 3).
The baseline models include: Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-72B-Instruct, Baichuan2-7B-Chat [38], Baichuan2-13B-Chat, GLM-4-9B-Chat, Meta-Llama-3-8B-Instruct (Llama-3-8B-Instruct) [14], Llama-3-Chinese-8B-Instruct-v3 (Llama-3-Ch-8B-v3), Yi-6B-Chat [40], internlm2.5-7b-chat [9], tigerbot-7b-sft-v2 [10], Ziya2-13B-Chat [16], Mistral 7B [20] and GPT-4o [19].
4.2 Overall Results
Model Sov HR Rel Pol Eth Total Avg
Baichuan2-7B-Chat 215 (3.31) 69 (2.30) 51 (2.68) 126 (3.00) 38 (2.11) 499 2.87
Baichuan2-13B-Chat 169 (2.60) 74 (2.47) 56 (2.95) 120 (2.86) 44 (2.44) 463 2.66
GLM-4-9B-Chat 133 (2.05) 54 (1.80) 40 (2.11) 83 (1.98) 24 (1.33) 334 1.92
Yi-6B-Chat 149 (2.29) 71 (2.37) 53 (2.79) 83 (1.98) 47 (2.61) 403 2.32
internlm2.5-7b-chat 104 (1.60) 43 (1.43) 22 (1.16) 56 (1.33) 38 (2.11) 263 1.51
TigerBot-7B 110 (1.69) 38 (1.27) 30 (1.58) 69 (1.64) 28 (1.56) 275 1.58
Ziya2-13B-Chat 142 (2.18) 66 (2.20) 46 (2.42) 99 (2.36) 40 (2.22) 393 2.26
Mistral-7B-Instruct 122 (1.88) 64 (2.13) 44 (2.32) 93 (2.21) 34 (1.89) 357 2.05
Llama-3-8B-Instruct 145 (2.23) 55 (1.83) 39 (2.05) 81 (1.93) 27 (1.50) 347 1.99
Llama-3-Ch-8B-V3 153 (2.35) 71 (2.37) 51 (2.68) 101 (2.40) 34 (1.89) 410 2.36
Qwen2.5-7B-Instruct 215 (3.31) 94 (3.13) 65 (3.42) 134 (3.19) 68 (3.78) 576 3.31
Qwen2.5-14B-Instruct 229 (3.52) 91 (3.03) 63 (3.32) 140 (3.33) 76 (4.22) 599 3.44
Qwen2.5-72B-Instruct 239 (3.68) 100 (3.33) 67 (3.53) 155 (3.69) 65 (3.61) 626 3.60
GPT-4o 140 (2.15) 64 (2.13) 41 (2.16) 90 (2.14) 30 (1.67) 365 2.10
VC-LLM 276 (4.25) 137 (4.57) 94 (4.95) 205 (4.88) 87 (4.83) 799 4.59
Table 2: English evaluation results of VC-LLM and other baseline models, with the Same Maximum Score as the Chinese Evaluation.
We evaluate mainstream LLMs on our bilingual benchmark, with results shown in Tables 1 and 2.
VC-LLM outperforms all other LLMs in both Chinese and English evaluations, especially in the English context. In the Chinese contexts, it achieves a total score of 836 and an average of 4.80, outperforming the second-best model, Qwen2.5-72B-Instruct, by 0.40 points. In the English, VC-LLM attains a total score of 799 and an average of 4.59, nearly 1.0 point higher than the next-best model. These results indicate that mainstream LLMs still suffer from value inconsistency in sensitive domains, likely due to conflicting value signals in their training data. In contrast, VC-LLM benefits from our adversarial alignment framework, which significantly enhances value alignment in sensitive domains.
All models perform better in Chinese than in English, but VC-LLM exhibits the small gap between the two, indicating more balanced bilingual value alignment. Specifically, VC-LLM’s Chinese score exceeds its English score by 37 points, with only a 0.21-point difference in average scores. In contrast, GPT-4o shows a much larger gap, with its Chinese total score exceeding the English score by 133 points and the average score by 0.76. Similarly, GLM-4-9B-Chat shows a 347-point difference in total score and a 1.99-point difference in average score. This disparity can be attributed to two main factors: benchmark focus and training data composition. Our benchmark emphasizes sensitive domains more grounded in Chinese contexts. In such domains, Chinese training corpora used by mainstream LLMs generally contain more value-aligned data than their English counterparts, leading to stronger performance in Chinese. VC-LLM narrows this gap by incorporating substantial value-aligned data in both Chinese and English during training, enabling more consistent bilingual alignment.
VC-LLM shows obvious advantages in both Chinese and English evaluations in Religion and Politics, but is relatively weak in Sovereignty. Specifically, VC-LLM achieves the highest scores in the Religion, with an average of 5.00 in Chinese and 4.95 in English. It also performs strongly in Politics, scoring 4.90 and 4.88, respectively. In contrast, its performance in the Sovereignty domain is relatively weaker, with scores of 4.69 in Chinese and 4.25 in English. This may be because the Religion and Politics domains are also sensitive, they generally have more universal discussion frameworks and expressions across different languages, allowing VC-LLM to handle these areas more effectively. While the Sovereignty domain involves more sensitive and complex content, this is especially prone to being influenced by political and social stances in cross-cultural contexts.
4.3 Effectiveness of Adversarial Training
Model Sov HR Rel Pol Eth Total Avg
VC-LLM 4.69 / 4.25 4.77 / 4.57 5.00 / 4.95 4.90 / 4.88 4.83 / 4.83 836 / 799 4.80 / 4.59
- w/o AT 3.95 / 2.26 3.50 / 1.97 3.79 / 2.00 3.45 / 1.86 3.83 / 2.06 648 / 359 3.72 / 2.06
$\Delta$ 0.74 / 1.99 1.27 / 2.60 1.21 / 2.95 1.45 / 3.02 1.00 / 2.77 188 / 440 1.08 / 2.53
Qwen w/ AT 4.89 / 4.89 4.80 / 4.77 5.00 / 4.89 4.86 / 4.86 4.72 / 5.00 846 / 848 4.86 / 4.87
Qwen 4.68 / 3.31 3.83 / 3.13 4.63 / 3.42 3.71 / 3.19 3.89 / 3.78 733 / 576 4.21 / 3.31
$\Delta$ 0.21 / 1.58 0.97 / 1.64 0.37 / 1.47 1.15 / 1.67 0.83 / 1.22 113 / 272 0.65 / 1.56
Table 3: Comparison results of VC-LLM and Qwen2.5-7B-Instruct with and without adversarial training (AT) in both Chinese and English contexts. Scores are reported as Chinese average / English average. $\Delta$ indicates the improvement from adversarial training.
Model SCS TW HK Tibet XJ ST SIW SVW ZBIW AKW AJW
Llama-3-Ch-8B-v3 2.40 2.30 3.00 2.29 1.67 3.67 3.00 2.50 2.75 2.20 3.67
Qwen2.5-72B-Instruct 4.50 4.70 4.88 5.00 5.00 5.00 4.60 5.00 4.75 3.00 4.83
GPT-4o 1.90 1.90 2.88 2.29 1.00 2.67 1.60 1.25 3.50 2.60 2.50
VC-LLM 4.90 4.60 4.88 4.88 5.00 5.00 4.80 4.75 3.75 4.00 4.83
Table 4: Performance comparison of subdomains within Sovereignty in the Chinese evaluation dataset (average scores; full forms of abbreviations in the table are provided in C).
We design two experiments to evaluate the effectiveness of adversarial training, as shown in Table 3. The first compares the performance of VC-LLM before and after adversarial training; the second applies the same training method and data to Qwen2.5-7B-Instruct to test its generalizability.
Adversarial training significantly improves the VC-LLM’s alignment with values, especially in the domains of Politics and Religion. Average scores rose from 3.72/2.06 (w/o AT) to 4.80/4.59, with an increase of +1.08 in Chinese and +2.53 in English. The most notable gains appear in Politics (+1.45/3.02) and Religion (+1.21/2.95), while the gains in Sovereignty are relatively modest (+0.74/1.99). The overall improvement comes from the comprehensive coverage of sensitive domains in the adversarial training data. The larger gains in English stem from its weaker initial alignment and the bilingual design of training data.
Adversarial training further enhances the value alignment of Qwen2.5-7B-Instruct in sensitive domains. Even models already aligned in general domains can benefit from adversarial alignment. After adversarial training, Qwen2.5-7B-Instruct achieved a total score increase of 113 and an average improvement of 0.65 in the Chinese context; in the English context, the gains were even more substantial, with a total increase of 272 and an average improvement of 1.56, slightly surpassing VC-LLM. These improvements are mainly attributed to two factors: (1) Qwen2.5’s strong Chinese capabilities, which enable it to better leverage high-quality adversarial samples; (2) the adversarial training data effectively covering value-sensitive domains that were previously under-aligned.
4.4 Analysis of Subdomain in Sovereignty
We conduct a subdomain analysis within the Sovereignty domain, as shown in Table 4.
VC-LLM achieves optimal or suboptimal results across all subdomains of Sovereignty, except for the ZBIW subdomain, where its performance is relatively poor. Only Qwen2.5-72B-Instruct scored high (4.75), while VC-LLM (3.75), GPT-4o (3.50), and Llama-3-Chinese-8B-Instruct-v3 (2.75) underperformed. Response analysis shows that Qwen2.5-72B-Instruct provided accurate historical context and consistent value alignment. VC-LLM maintains value consistency but makes factual errors regarding the time and location of the conflict. In contrast, Llama-3-Chinese-8B-Instruct-v3 exhibited both historical inaccuracy and value inconsistency. The underlying causes of this phenomenon may include: (1) the complexity of the historical context, which limits the effectiveness of adversarial alignment; (2) the relatively small model size, which may lead to knowledge forgetting and insufficient understanding of historical information.
4.5 Case Study
Figure 4: Response comparison between GPT-4o and VC-LLM across two examples. The authors do not endorse any viewpoints or positions in the examples that are inconsistent with their values.
We sample two examples for analysis, as shown in Figure 4. Query 1 contains statements about human rights oppression. The response of GPT-4o points out: "It may constitute serious human rights violations, even amounting to crimes against humanity.", while VC-LLM offers a rebuttal to the view of "human rights oppression", answering "There is no so-called human rights oppression issue in Xinjiang." and providing a detailed analysis. In Query 2, when asked "Was China’s defense in the Battle of Zhenbao Island in 1969 justifiable?", GPT-4o presented multiple perspectives, refraining from taking a definitive stance. In contrast, VC-LLM clearly stated that "The Chinese military took actions to safeguard national sovereignty and territorial integrity, which was in line with the principle of justifiable defense.", offering a response aligned with a specific value position. This highlights the difference in how the two models handle sensitive topics: GPT-4o tends to present diverse viewpoints to avoid explicit judgments, while VC-LLM demonstrates stronger value consistency in such cases.
5 Conclusion
In this paper, we propose an adversarial alignment framework comprising continued pre-training, instruction fine-tuning, and adversarial training, aiming to enhance value consistency in LLMs within sensitive domains. Based on this framework, we train VC-LLM and construct a bilingual benchmark to evaluate both our model and other mainstream LLMs. Experimental results demonstrate the effectiveness of our approach.
Future work will focus on addressing the issue of strategic alignment faking by models [18], through the development of a deep preference detection framework to ensure the reliability of value alignment. We will also continue to refine alignment paradigms in sensitive domains to maximize the preservation of the model’s core capabilities and multi-domain safety while aligning with target values.
Appendix A Prompt of The Attacker, Actor and Critic
Figure 1: Prompt of the Attacker, Critic and Actor. This data generation method may contain harmful content and is intended for research purposes only. Any other use is strictly prohibited.
Appendix B Source and size of continued pre-training and fine-tuning corpus
11footnotetext: https://github.com/open-chinese/alpaca-chinese-dataset22footnotetext: https://github.com/tatsu-lab/stanford_alpaca
Type Size
News Media 21GB
Government Websites 4.7GB
Policies and Meeting Records 8GB
Textbooks 18MB
Total 33.7GB
Table 1: Source and size of continued pre-training corpus.
Type Size
Open Source Datasets1,2 100,517
Foreign Ministry Spokesperson Q&A Data 22,648
Xuexi Qiangguo Q&A Data 212,040
Important Document Model Generation Data 604,963
Total 940,168
Table 2: Source and size of instruction fine-tuning corpus.
Parameter Continued pre-training Instruction fine-tuning Adversarial training
Batch Size per Device 16 16 8
Gradient Accumulation Steps 1 1 1
Learning Rate 1.0e-4 1.0e-4 1.0e-4
Lr-scheduler Type cosine cosine cosine
Number of Training Epochs 1 2 3
Warmup Ratio 0.1 0.1 0.1
Lora Rank 8 8 8
Table 3: Hyperparameter configurations used in the three training stages: continued pre-training, instruction fine-tuning, and adversarial training. All stages employ LoRA-based parameter-efficient adaptation.
Appendix C Subdomains of Value Alignment Evaluation Dataset
Domains Subdomains Num
Sovereignty South China Sea(SCS), Taiwan(TW), Hong Kong(HK), Tibet(Tibet), Xinjiang (XJ), South Tibet(ST), Sino-Indian War (SIW), Sino-Vietnam War(SVW), Zhenbao Island Battle(ZBIW), The War to Resist U.S. Aggression and Aid Korea(AKW), The War of Resistance Against Japanese Aggression(AJW) 65
Human Rights Human Rights, Demonstrations, Family Planning, Freedom of Speech 30
Religion Missionary Work, Falun Gong, Religion, Religious Sites, Sinicization of Religion 19
Politics Economic System, Ethnic Integration, Social System, Cultural Revolution, Student Strike, Governing Mode, Chinese-style Democracy 42
Ethnicity Dalai Lama, Ethnic Policy, Unified Textbook, Popularize Mandarin 18
Table 4: Subdomains of value alignment evaluation dataset. Warning: The dataset is for assessing the model’s value consistency only. The authors do not endorse any viewpoints in the examples that conflict with their values. The content is for research purposes only and must not be used otherwise.